# Practical learning of multi-time statistics in open quantum systems

### v1

@@ 1 Introduction
Specifically, in this work, we address the question: how can many-time phenomena, as described above, be accessed in quantum experiments and on near-term quantum hardware? In doing so, we pave the way for the realistic observation and study of temporal correlations, placing them on a similar level as many-body states. This has timely applications across quantum technologies, including model-independent environmental characterization, novel sensing methods, diagnosis of temporally correlated noise, and paths to quantum advantage through the simulation of open systems. To achieve this, we combine established approaches in many-body physics with new techniques specifically designed for observing temporal quantum states, particularly under various control paradigms. We show that quantum computers, even in their current state, offer a timely opportunity to explore the richness of many-time physics. While these machines are too noisy to manipulate into the large entangled registers required for interesting quantum algorithms, they are riddled with complex noise [26, 23, 25], which we can access with high-fidelity controls. Thus, we show that there is scope to explore interesting physics even with relatively small devices – a single qubit probed across many times, for example, can exhibit many of the same properties as large quantum states. In this sense, these methodological advances show that highly nontrivial many-time physics can be more accessible than many-body physics for the near-term quantum processors, which opens rich new areas of investigation. Applications of this work range from the diagnostics of temporally correlated noise in quantum technologies [27, 26], to the observation and simulation of dynamical complexity in condensed matter systems [28, 29, 30], to quantum biology [31, 32] and quantum causal modelling [3].
@@ 2.1 Quantum channels
^$\hat{\mathcal{E}}:=(\mathcal{E}\otimes\mathcal{I})\left[|\Phi^{+}\rangle\langle\Phi^{+}|\right]=\sum_{i,j=1}^{d}\mathcal{E}\left[|i\rangle\langle j|\right]\otimes|i\rangle\langle j|.$ (3)
@@
In the lab, one can characterise the map $\mathcal{E}$ in Eq. (1) by a suitable choice of preparations and measurements. However, one does not have access to the environment and the unitary quantum transformations in Eq. (2). Nevertheless, the Stinespring dilation theorem [7] guarantees that for every completely positive map, there exists a system-environment unitary transformation of the form of the right-hand side of the same equation. Note that this is only under the condition that initial system-environment correlations are not present [Pechukas1994].
@@ 2.1.1 Instruments
Figure 2: a We refer to the Choi states of quantum processes as many-time states, which are isomorphic to many-body states modulo causality conditions. Hence, we can employ many of the tools developed to study many-body states to study quantum stochastic processes. b Many-time non-Markovian processes are represented by process tensors, which is a multitime density matrix that contracts with a sequence of instruments to yield a final state for the system. c A $k$-step process tensor can also be seen as a sequence of correlated CPTP maps between different points in time, plus the effect due to the initial state. Output legs $\mathfrak{o}_{l}$ are mapped by a control operation $\mathcal{A}_{l}$ to the next input leg $\mathfrak{i}_{l+1}$.
@@ 2.3 Properties of quantum processes
where {^$\{\hat{\mathcal{E}}_{y:x}^{\mu}\}$are CP maps from time $x$ to $y$. The index $\mu$ and the amplitudes $\alpha_{\mu}$ account for correlations in time. This is depicted in Figure 2b. For classical non-Markovian processes, each ^$\hat{\mathcal{E}}^{\mu}$ is a CPTP channel and $\alpha_{\mu}\geq 0$.
@@ 2.4 Process tensor tomography
The final instrument becomes a POVM (positive operator-valued measure) because we do not require the post-measurement state at that stage. [^4] If the probabilities are known only for a subset of operations (in this work, we are often restricted to the space of unitary operations), the process tensor will be uniquely determined within that subspace [25, 39].
@@ 3 Spatiotemporal Classical Shadows
In this section, we will extend the state-of-the-art in learning properties of quantum states to quantum stochastic processes. To do so, we will make use of the properties of quantum stochastic processes discussed in the last section, both theoretical and practical perspectives. This extension will enable learning the properties of non-Markovian quantum stochastic processes, even when dealing with large numbers of steps or qubits.
@@ 3.2 Temporal Classical Shadows
At each time $t_{j}$, the process has an output leg $\mathfrak{o}_{j}$ (which is measured), and input leg $\mathfrak{i}_{j+1}$ (which is a state that feeds back into the process). The two important properties that we stress are: (i) a sequence of operations constitutes an observable on the process tensor via Eq. (14), generating the connection to classical shadows, and (ii) a process tensor constitutes a collection of possibly correlated CPTP maps and hence may be marginalised in both time and space to yield the $j$th CPTP map describing the dynamics of the $i$th qubit ^$\hat{\mathcal{E}}^{(q_{i})}_{j:j-1}$.
@@
In the tomographic representation of a process tensor (see Ref. [2]), given a basis of operations at the $i$th time $\{\mathcal{B}_{i}^{\mu_{i}}\}$ with Choi states {^$\{\hat{\mathcal{B}}_{i}^{\mu_{i}}\}$. There exists a dual set $\mathcal{D}:=\{\Delta_{j}^{\mu_{j}}\}$ such that the process tensor may be written
@@
To compute the classical snapshot from a sequence of instruments, first, let $N$ be the size of the (possibly under or overcomplete) basis. We can write each of the Choi ^$\hat{\mathcal{B}}_{i}^{\mu_{i}}$ as a row vector (adopting a row-vectorised convention) ^$\langle\!\langle\hat{\mathcal{B}}_{i}^{\mu_{i}}|$ and from this construct a single matrix ^$\hat{\mathcal{M}}^{\prime}_{i}$:
^$\hat{\mathcal{M}}^{\prime}_{i}=\begin{pmatrix}\langle\!\langle\hat{\mathcal{B}}_{i}^{0}|\\ \langle\!\langle\hat{\mathcal{B}}_{i}^{1}|\\ \vdots\\ \langle\!\langle\hat{\mathcal{B}}_{i}^{N-1}|\\ \end{pmatrix}.$ (25)
Now, the matrix $\mathcal{M}_{i}^{\prime\ +\dagger}$ (right-inverse, conjugate transposed) contains the dual matrices $\{\Delta_{i}^{\mu_{i}}\}$ to {^$\{\hat{\mathcal{B}}_{i}^{\mu_{i}}\}$ in its rows, which are then trace normalised.
@@ 3.3.1 Low Weight Observables
The implication here is that for any feasible set of experiments, the desired operators to learn must be mostly $\mathbb{I}$, with a handful of $X$, $Y$, and $Z$. We recall that $\mathbb{I}$ in the Choi picture is either a partial trace (measure and discard) in the outputs, or a preparation of a maximally mixed state in the inputs. In quantum states, there is no a priori structure to the set of correlations between qubits. However, with processes, we have seen in Sec. 2 the effects of causality constraints. These ensure that there is a causal direction. But this is problematic, because information must travel in and out of the system – and so any discarding of the system by low-weight observables equally erases important information. One concrete example is in a two-step unital quantum process from which we wish to extract two-point correlation functions. Causality forbids any $\mathfrak{i}_{1}-\mathfrak{i}_{2}$, and $\mathfrak{i}_{1}-\mathfrak{o}_{0}$ or $\mathfrak{i}_{2}-\mathfrak{o}_{1}$ correlations; unitality forbids $\mathfrak{o}_{2}-\mathfrak{o}_{1}$, $\mathfrak{o}_{1}-\mathfrak{o}_{0}$, and $\mathfrak{o}_{2}-\mathfrak{o}_{0}$ correlations. This leaves $\mathfrak{i}_{1}-\mathfrak{o}_{2}$ as the only non-trivial two-point correlator of the process.
@@ 3.3.2 Process Marginals
We find that one particularly useful application of classical shadows to quantum stochastic processes is simultaneous determination of process tensor marginals, particularly contiguous ones. This allows us to determine a subset of correlations between different times. More usefully, though, we shall explore different classes of processes that may be reconstructed from their marginals. A $k-$step process marginal across $n$ qubits is fully determined by a set of $\mathcal{O}(2^{2kn})$ weight $kn$ Pauli observables. When marginalising across all but a handful of times or qubits, we will denote the remaining steps or registers by commas, i.e.
^$\begin{split}\hat{\mathcal{E}}^{(q_{i_{0}},q_{i_{1}})}_{j_{0}:j_{0}-1,j_{1}:j_{1}-1}\!&:=\!\text{Tr}_{{\{\overline{q_{i_{0}}},\overline{q_{i_{1}}}\}},\{\overline{t_{j_{0}}},\overline{t_{j_{0}-1}},\overline{t_{j_{1}}},\overline{t_{j_{1}-1}}\}}[\Upsilon_{k:0}];\\ \Upsilon_{k:0}^{(q_{i})}&=\text{Tr}_{\overline{q_{i}}}[\Upsilon_{k:0}],\end{split}$ (30)
where the overlines denote complement. For example, if $\mathbf{Q}:=\{q_{0},q_{1},\cdots,q_{n-1}\}$ were to denote all qubits on a device, then $\overline{q_{i}}=\mathbf{Q}\backslash\{q_{i}\}$. However, for the remainder of this work we operate only on a single qubit and omit this label.
Although collecting shadows suffices to estimate properties of the process marginals, they will not in general be able to reconstruct a physical estimate of the process marginals. To this effect, we estimate enough observables to constitute informationally complete information about the process tensor and then employ MLE-PTT to process the data and obtain a physical estimate. The advantage of a physical (positive, causal) estimate is that we make use of information-theoretic tools that rely on the positivity of the state.
@@ 4 Manufacturing Informationally Complete Quantum Instruments
Below we will make use of the standard one-qubit quantum logic gates. Let $H$ denote the Hadamard gate, $F$ denote the phase gate, and $X,\ Y,\ Z$ denote the Pauli gates. Each of these is defined as:
@@ 4.1 Engineering Informationally-Complete Control
where $V_{\gamma}$ is a cross-resonance $SA$ interaction [50] that leads to a short non-Markovian process as per Eq. (4). This is the native cross-resonance interaction on IBM Quantum devices – for some $\gamma$ [^6] The cross-resonance interactions are likely introduce extraneous local terms, so what results will not be pure $XZ$ but will be characterised nonetheless. Next, a phase gate $F$ is then applied on $A$ and a unitary $w$ on $S$ parametrised as
@@
Figure 3: Instrument design for complete multi-time statistics. a Using a structured system-ancilla interaction, we construct a restricted process tensor with local unitaries to realise the desired mapping. This gives an exact characterisation of the effective instrument $\mathcal{A}_{x}(\theta,\phi,\lambda)$ for any choice of parameter values. b Some sample basis instruments implemented using this method on ibm_perth written in the Pauli basis, showing IC controllability including non-unital and trace-decreasing directions. c Many-time physics is then studied using the circuit structured as shown. The different instruments applied capture temporal quantum correlations from an arbitrarily strong $SE$ interaction.
@@ 5 Correlated Noise Characterisation
We demonstrate this approach on an IBM Quantum device, obtaining these detailed diagnostics across many different time steps from only a relatively small number of shots.In this work, we perform experiments on devices with mid-circuit measurement capabilities. Note that the choice to use an ancilla is because these measurements typically have a long duration (1–5 $\mu$s). Deferring to an ancilla allows the system to continue participating in dynamics – otherwise, the window is lost if measuring the system directly. We employ the bootstrapped IC basis using the procedure outlined in Sec. 4, summarised in Fig. 3. For the local unitary operation, random Clifford operations are applied to the system qubit between system-ancilla interactions. The circuits are therefore of the form given in Fig. 3c, and for each of the $N$ outcomes, a snapshot of the form of Eq. 26 constructed to form a classical shadow $\mathtt{S}(\Upsilon_{k:0},N)$.
@@ 5.1 Correlated errors
^$\hat{\mathcal{E}}_{j+1:j;i+1:i}:=\text{Tr}_{\bar{i},\bar{j}}[\Upsilon_{k:0}],$ (39)
@@
We collected $10^{7}$ shots of randomised instruments and then use Eq. (26) to construct each of the corresponding classical shadows. The data is then further postprocessed to determine IC information about the marginals, and an MLE estimate is obtained. These estimates give us several representative components of the noise. Specifically, Figure 4 displays the negativity between each pair of maps; the trace distance between the correlated pairs and tensor product of marginals; and the overlap of these maps with Pauli error channels. Note that for marginals between non-neighbouring times, there is a maximal depolarising channel applied in the interim. However, we see more sensitivity here to the exact multi-time correlations, owing to the use of an IC probe to distinguish different aspects of the process. While we only considered a seven-steps process, the procedure scales only logarithmically with the number of times and so in principle could be extended far further.
Figure 4: The two-map (four-body) marginals of a system enduring naturally idle dynamics on ibm_perth. a We compute both the negativity between the two steps and the trace-distance between the maps and the product of their marginals. b We compute the joint overlap of correlated Pauli errors (their Choi state): $\langle\hat{P}_{i}\hat{P}_{j}\rangle-\langle\hat{P}_{i}\rangle\langle\hat{P}_{j}\rangle$ for $ZZ$ and $XX$ as well as c $ZX$ and $XZ$.

### v2

@@ 1 Introduction
Specifically, in this work, we address the question: how can many-time phenomena, as described above, be accessed in quantum experiments and on near-term quantum hardware? In doing so, we pave the way for the realistic observation and study of temporal correlations, placing them on a similar level as many-body states. To achieve this, we combine established approaches in many-body physics with new techniques specifically designed for observing temporal quantum states, particularly under various control paradigms. We show that quantum computers, even in their current state, offer a timely opportunity to explore the richness of many-time physics. While these machines are too noisy to manipulate into the large entangled registers required for interesting quantum algorithms, they are riddled with complex noise [26, 23, 25], which we can access with high-fidelity controls. Thus, we show that there is scope to explore interesting physics even with relatively small devices – a single qubit probed across many times, for example, can exhibit many of the same properties as large quantum states. In this sense, these methodological advances show that highly nontrivial many-time physics can be more accessible than many-body physics for the near-term quantum processors, which opens rich new areas of investigation. Applications of this work range from the diagnostics of temporally correlated noise in quantum technologies [27, 26], to the observation and simulation of dynamical complexity in condensed matter systems [28, 29, 30], to quantum biology [31, 32] and quantum causal modelling [3].
@@ 2.1 Quantum channels
$\hat{\mathcal{E}}:=(\mathcal{E}\otimes\mathcal{I})\left[|\Phi^{+}\rangle\langle\Phi^{+}|\right]=\sum_{i,j=1}^{d}\mathcal{E}\left[|i\rangle\langle j|\right]\otimes|i\rangle\langle j|.$ (3)
@@
In the lab, one can characterise the map $\mathcal{E}$ in Eq. (1) by a suitable choice of preparations and measurements. However, one does not have access to the environment and the unitary quantum transformations in Eq. (2). Nevertheless, the Stinespring dilation theorem [7] guarantees that for every completely positive map, there exists a system-environment unitary transformation of the form of the right-hand side of the same equation. Note that this is only under the condition that initial system-environment correlations are not present [33].
@@ 2.1.1 Instruments
Figure 2: a We refer to the Choi states of quantum processes as many-time states, which are isomorphic to many-body states modulo causality conditions. Hence, we can employ many of the tools developed to study many-body states to study quantum stochastic processes. b A many-time non-Markovian process is represented by a process tensor, which is a multitime density matrix that contracts with a sequence of instruments to yield a final state for the system. c A $k$-step process tensor can also be seen as a sequence of correlated CPTP maps between different points in time, plus the effect due to the initial state. Output legs $\mathfrak{o}_{l}$ are mapped by a control operation $\mathcal{A}_{l}$ to the next input leg $\mathfrak{i}_{l+1}$.
@@ 2.2 Process tensors
We emphasise our conventions on time throughout this work: in circuit diagrams, time runs from left to right in accordance with the literature. In matrix and tensor representations, however, time runs from right to left, in accordance with the standards of matrix multiplication. Although the tension between these two is unfortunate, we believe it simplifies matters in practice.
@@ 2.3 Properties of quantum processes
where $\{\hat{\mathcal{E}}_{y:x}^{\mu}\}$are CP maps from time $x$ to $y$. The index $\mu$ and the amplitudes $\alpha_{\mu}$ account for correlations in time. This is depicted in Figure 2b. For classical non-Markovian processes, each $\hat{\mathcal{E}}^{\mu}$ is a CPTP channel and $\alpha_{\mu}\geq 0$.
@@ 2.4 Process tensor tomography
The final instrument becomes a POVM (positive operator-valued measure) because we do not require the post-measurement state at that stage. [^4] If the probabilities are known only for a subset of operations.
@@ 3 Spatiotemporal Classical Shadows
In this section, we will extend the state-of-the-art in learning properties of quantum states to quantum stochastic processes. To do so, we will make use of the properties of quantum stochastic processes discussed in the last section, from both theoretical and practical perspectives. This extension will enable learning the properties of non-Markovian quantum stochastic processes, even when dealing with large numbers of steps or qubits.
@@ 3.2 Temporal Classical Shadows
At each time $t_{j}$, the process has an output leg $\mathfrak{o}_{j}$ (which is measured), and input leg $\mathfrak{i}_{j+1}$ (which is a state that feeds back into the process). The two important properties that we stress are: (i) a sequence of operations constitutes an observable on the process tensor via Eq. (14), generating the connection to classical shadows, and (ii) a process tensor constitutes a collection of possibly correlated CPTP maps and hence may be marginalised in both time and space to yield the $j$th CPTP map describing the dynamics of the $i$th qubit $\hat{\mathcal{E}}^{(q_{i})}_{j:j-1}$.
@@
In the tomographic representation of a process tensor (see Ref. [2]), given a basis of operations at the $i$th time $\{\mathcal{B}_{i}^{\mu_{i}}\}$ with Choi states $\{\hat{\mathcal{B}}_{i}^{\mu_{i}}\}$. There exists a dual set $\mathcal{D}:=\{\Delta_{j}^{\mu_{j}}\}$ such that the process tensor may be written
@@
To compute the classical snapshot from a sequence of instruments, first, let $N$ be the size of the (possibly under or overcomplete) basis. We can write each of the Choi $\hat{\mathcal{B}}_{i}^{\mu_{i}}$ as a row vector (adopting a row-vectorised convention) $\langle\!\langle\hat{\mathcal{B}}_{i}^{\mu_{i}}|$ and from this construct a single matrix $\hat{\mathcal{M}}^{\prime}_{i}$:
$\hat{\mathcal{M}}^{\prime}_{i}=\begin{pmatrix}\langle\!\langle\hat{\mathcal{B}}_{i}^{0}|\\ \langle\!\langle\hat{\mathcal{B}}_{i}^{1}|\\ \vdots\\ \langle\!\langle\hat{\mathcal{B}}_{i}^{N-1}|\\ \end{pmatrix}$. (25)
Now, the matrix $\mathcal{M}_{i}^{\prime\ +\dagger}$ (right-inverse, conjugate transposed) contains the dual matrices $\{\Delta_{i}^{\mu_{i}}\}$ to $\{\hat{\mathcal{B}}_{i}^{\mu_{i}}\}$ in its rows, which are then trace normalised.
@@ 3.3.1 Low Weight Observables
The implication here is that for any feasible set of experiments, the desired operators to learn must be mostly $\mathbb{I}$, with a handful of $X$, $Y$, and $Z$. In the Choi picture, the operator $\mathbb{I}$ has an operational meaning as the physical instrument of either a partial trace (measure and discard) in the outputs $\mathfrak{o}_{j}$, or a preparation of a maximally mixed state in the inputs $\mathfrak{i}_{j}$. As an instrument across $\mathfrak{o}_{j-1}\mathfrak{i}_{j}$, it represents a completely depolarising channel. In quantum states, there is no a priori structure to the set of correlations between qubits. However, with processes, we have seen in Sec. 2 the effects of causality constraints. These ensure that there is a causal direction. But this is problematic, because information must travel in and out of the system – and so any discarding of the system by low-weight observables equally erases important information. One concrete example is in a two-step unital quantum process from which we wish to extract two-point correlation functions. Causality forbids any $\mathfrak{i}_{1}-\mathfrak{i}_{2}$, and $\mathfrak{i}_{1}-\mathfrak{o}_{0}$ or $\mathfrak{i}_{2}-\mathfrak{o}_{1}$ correlations; unitality forbids $\mathfrak{o}_{2}-\mathfrak{o}_{1}$, $\mathfrak{o}_{1}-\mathfrak{o}_{0}$, and $\mathfrak{o}_{2}-\mathfrak{o}_{0}$ correlations. This leaves $\mathfrak{i}_{1}-\mathfrak{o}_{2}$ as the only non-trivial two-point correlator of the process.
@@ 3.3.2 Process Marginals
We find that one particularly useful application of classical shadows to quantum stochastic processes is simultaneous determination of process tensor marginals, particularly contiguous ones. This allows us to determine a subset of correlations between different times. More usefully, though, we shall explore in Sec. 5.2 different classes of processes that may be reconstructed from their marginals. A $k$-step process marginal across $n$ qubits is fully determined by a set of $\mathcal{O}(2^{2kn})$ weight $kn$ Pauli observables. When marginalising across all but a handful of times or qubits, we will denote the remaining steps or registers by commas, i.e.
$\begin{split}\hat{\mathcal{E}}^{(q_{i_{0}},q_{i_{1}})}_{j_{0}:j_{0}-1,j_{1}:j_{1}-1}\!&:=\!\text{Tr}_{{\{\overline{q_{i_{0}}},\overline{q_{i_{1}}}\}},\{\overline{t_{j_{0}}},\overline{t_{j_{0}-1}},\overline{t_{j_{1}}},\overline{t_{j_{1}-1}}\}}[\Upsilon_{k:0}];\\ \Upsilon_{k:0}^{(q_{i})}&=\text{Tr}_{\overline{q_{i}}}[\Upsilon_{k:0}],\end{split}$ (30)
[…]
Although collecting shadows suffices to estimate properties of the process marginals, they will not in general be able to reconstruct a physical estimate of the process marginals themselves. To this effect, we estimate enough observables to constitute informationally complete information about the process tensor and then employ MLE-PTT to process the data and obtain a physical estimate. The advantage of a physical (positive, causal) estimate is that we make use of information-theoretic tools that rely on the positivity of the state.
Lastly, it is important to ascribe and emphasise an appropriate physical significance to process marginals. In a multi-partite state $\rho_{ABC}$, the state marginal $\rho_{AC}$ (i.e., after performing a partial trace on $B$) is the state one obtains after projectively measuring $B$ and discarding the outcome. In a multi-partite channel $\mathcal{E}_{ABC}$, the channel marginal $\mathcal{E}_{AC}$ is the resultant channel one would obtain by conditionally inserting a maximally mixed state on system $B$, and then projectively measuring $B$ and discarding the outcome. Process marginals inherit this same interpretation, except that now the multi-party channel corresponds to different inputs and outputs on the same system at various times. That is, for example, the process marginal $\hat{\mathcal{E}}_{3:2;1:0}$ is the conditional process one would obtain by inserting a maximally mixed state at time $t_{1}$ and measuring/discarding at $t_{2}$. Although this might seem like a distant abstraction from physical settings, it essentially corresponds to being ignorant of the system for a given window of time. For example, the single-step marginal $\hat{\mathcal{E}}_{j}$ is the resulting channel one would obtain if they performed quantum process tomography only for the evolution from time $t_{j}$ to $t_{j+1}$.
@@ 4 Manufacturing Informationally Complete Quantum Instruments
Below we will make use of the standard one-qubit quantum logic gates. Let $H$ denote the Hadamard gate, $F$ denote the phase gate, and X,Y,ZX,\ Y,\ Z denote the Pauli gates. Each of these is defined as:
@@ 4.1 Engineering Informationally-Complete Control
where $V_{\gamma}$ is a cross-resonance $SA$ interaction [51] that leads to a short non-Markovian process as per Eq. (4). This is the native cross-resonance interaction on IBM Quantum devices – for some $\gamma$. [^6] The cross-resonance interactions are likely introduce extraneous local terms, so what results will not be pure $XZ$ but will be characterised nonetheless. Next, a phase gate $F$ is then applied on $A$ and a unitary $w$ on $S$ parametrised as
@@
Figure 3: Instrument design for complete multi-time statistics. a Using a structured system-ancilla interaction, we construct a restricted process tensor with local unitaries to realise a locally-controllable instrument. This gives an exact characterisation of the effective instrument $\mathcal{A}_{x}(\theta,\phi,\lambda)$ for any choice of parameter values. Importantly, as per Eq. (37), these instruments will access the full trace-decreasing and non-unital subspace of superoperator space. b Some sample basis instruments implemented using this method on ibm_perth written in the Pauli basis, showing IC controllability including non-unital and trace-decreasing directions. These PTMs demonstrate that well-conditioned IC instrument sets can be achieved.c An illustration of how these instrument sets can be used to probe multi-time statistics. Each multi-time instrument is realised with the circuit as shown – by applying the locally-modulated interaction with the ancilla, measuring (and recording) the ancilla, and subsequently resetting it. The different instruments applied capture temporal quantum correlations from an arbitrarily strong SESE interaction.
@@ 5 Correlated Noise Characterisation
[…]
@@ 5.1 Correlated errors
$\hat{\mathcal{E}}_{j+1:j;i+1:i}:=\text{Tr}_{\bar{i},\bar{j}}[\Upsilon_{k:0}],$ (39)
@@
We collected $10^{7}$ shots of randomised instruments and then use Eq. (26) to construct each of the corresponding classical shadows. The data is then further postprocessed to determine IC information about the marginals, and an MLE estimate is obtained. These estimates give us several representative components of the noise. Specifically, Figure 4 displays:
  • The negativity between each pair of maps. That is, $\mathcal{N}_{j\mid i}[\hat{\mathcal{E}}_{j+1:j;i+1:i}]$, where $\mathcal{N}_{A\mid B}[\rho]:=\frac{1}{2}(\|\rho^{\Gamma_{A}}\|_{1}-1)$ is an entanglement monotone ($\Gamma_{A}$ is the partial transpose with respect to $A$), and can only be non-zero for genuinely quantum memory [55].
  • The trace distance $\frac{1}{2}\|\hat{\mathcal{E}}_{j+1:j;i+1:i}-\hat{\mathcal{E}}_{j+1:j}\otimes\hat{\mathcal{E}}_{i+1:i}\|_{1}$, measuring distinguishability of the joint two steps from the tensor product of their marginals (which would be obtained, e.g., from quantum process tomography experiments).
  • The connected Pauli probabilities across the two steps. That is, defining $\langle\hat{P}_{i}\hat{P}_{j}\rangle$ to be $\text{Tr}[\hat{\mathcal{E}}_{j+1:j;i+1:i}(\hat{P}_{j}\otimes\hat{P}_{i})]$, where $\hat{P}$ is the Choi state of a Pauli operator, then we plot $\langle\hat{P}_{i}\hat{P}_{j}\rangle-\langle\hat{P}_{i}\rangle\langle\hat{P}_{j}\rangle$.
Note that for marginals between non-neighbouring times, there is a maximal depolarising channel applied in the interim. However, we see more sensitivity here to the exact multi-time correlations, owing to the use of an IC probe to distinguish different aspects of the process. While we only considered a seven-steps process, the procedure scales only logarithmically with the number of times and so in principle could be extended far further.
Figure 4: The two-map (four-body) marginals of a system enduring naturally idle dynamics on ibm_perth. a We compute both the negativity between the two steps and the trace-distance between the maps and the product of their marginals. b We compute the joint overlap of correlated Pauli errors (their Choi states): $\langle\hat{P}_{i}\hat{P}_{j}\rangle-\langle\hat{P}_{i}\rangle\langle\hat{P}_{j}\rangle$ for ZZZZ and XXXX as well as c ZXZX and $XZ$.